# Constraints on Genesis Cosmology from the Smeared Null Energy Condition

### v1

@@ I Introduction
Nonsingular cosmic bounce Gasperini:1992em; Finelli:2001sr; Piao:2003zm; Cai:2007qw; Easson:2011zy; Qiu:2011cy; Cai:2012va; Liu:2013kea; Koehn:2013upa; Qiu:2013eoa; Nojiri:2016ygo; Li:2016awk; Cai:2017pga; Ye:2019sth; Ilyas:2020qja; Zhu:2021whu; Zhu:2023lhv and Genesis Creminelli:2010ba; Liu:2011ns; Wang:2012bq; Liu:2012ww; Creminelli:2012my; Hinterbichler:2012fr; Hinterbichler:2012yn; Easson:2013bda; Liu:2014tda; Pirtskhalava:2014esa; Nishi:2015pta; Cai:2016gjd; Nishi:2016ljg; Mironov:2019qjt; Ilyas:2020zcb; Zhu:2021ggm; Cai:2022ori are two representative scenarios for addressing the past geodesic incompleteness of inflation, and both involve violating the NEC. Theoretical studies in general relativity indicate that the geometric structure of a traversable wormhole requires matter fields that violate the NEC to sustain the throat Morris:1988cz; Morris:1988tu. Recently, observations of baryon acoustic oscillations by the DESI collaboration DESI:2025fii have also sparked interest in whether dark energy in the late-time universe violates the NEC Ye:2025ulq; Moghtaderi:2025cns; Caldwell:2025inn. In addition, NEC violation also provides a compelling mechanism for the formation of PBHs as a dark matter candidate Cai:2023uhc, as well as for generating stochastic gravitational-wave background signals detectable by pulsar timing array observations Zhu:2023lbf; Ye:2023tpz, see also Cai:2020qpu; Cai:2022nqv; Cai:2022lec; Pan:2024ydt; Li:2024oru; Li:2025ilc; Lai:2025xov; Lai:2025efh.
Although widely employed in theoretical models, the viability of NEC violation in the real world remains in question. Under the perfect-fluid assumption, NEC violation means the total energy density plus pressure becomes negative along a null direction, which allows exotic behaviors like an increase of energy density even in an adiabatically expanding universe. For theorists, NEC violation is often viewed unfavorably because it tends to trigger pathological instabilities, including ghost and gradient instabilities Libanov:2016kfc; Kobayashi:2016xpl, see also Battarra:2014tga; Koehn:2015vvy; Ijjas:2016tpn. It is demonstrated in the context of nonsingular cosmology that fully stable NEC violation can be constructed in “beyond Horndeski” theories Cai:2016thi; Creminelli:2016zwa; Cai:2017tku; Cai:2017dyi; Kolevatov:2017voe. Furthermore, the NEC can notoriously be violated at the quantum level, for instance in the quantum Casimir effect Brown:1969na and in the Hawking evaporation process (see, e.g. Kontou:2023ntd for a recent update on this topic).
@@ II.2 The SNEC conjecture
where $t_{\mathrm{s}}$ and $t_{\mathrm{e}}$ denote the beginning and the end of the cosmic interval we are interested in, respectively. Here, we have already used $\mathrm{d}\lambda=a\mathrm{d}t$. In the following, (10) serves as the master condition for examining the Genesis scenario.
@@ Acknowledgements.
We thank Yun-Song Piao for valuable discussions. Dong-Hui Yu and Yong Cai is supported in part by the National Natural Science Foundation of China (Grant Nos. 12575066, 11905224) and the Natural Science Foundation of Henan Province (Grant No. 242300420231). Mian Zhu is supported by the National Natural Science Foundation of China (Grant No. 12503005), and the Fundamental Research Funds for the Central Universities Grant No. YJ202551.

### v2

@@ I Introduction
Nonsingular cosmic bounce Gasperini:1992em ; Finelli:2001sr ; Piao:2003zm ; Cai:2007qw ; Easson:2011zy ; Qiu:2011cy ; Cai:2012va ; Liu:2013kea ; Koehn:2013upa ; Qiu:2013eoa ; Nojiri:2016ygo ; Li:2016awk ; Cai:2017pga ; Ye:2019sth ; Battista:2020lqv ; Ilyas:2020qja ; Battista:2022hqn ; Zhu:2021whu ; Zhu:2023lhv and Genesis Creminelli:2010ba ; Liu:2011ns ; Wang:2012bq ; Liu:2012ww ; Creminelli:2012my ; Hinterbichler:2012fr ; Hinterbichler:2012yn ; Easson:2013bda ; Liu:2014tda ; Pirtskhalava:2014esa ; Nishi:2015pta ; Cai:2016gjd ; Nishi:2016ljg ; Mironov:2019qjt ; Ilyas:2020zcb ; Zhu:2021ggm ; Cai:2022ori are two representative scenarios for addressing the past geodesic incompleteness of inflation, and both involve violating the NEC. Theoretical studies in general relativity indicate that the geometric structure of a traversable wormhole requires matter fields that violate the NEC to sustain the throat Morris:1988cz ; Morris:1988tu . Recently, observations of baryon acoustic oscillations by the DESI collaboration DESI:2025fii have also sparked interest in whether dark energy in the late-time universe violates the NEC Ye:2025ulq ; Moghtaderi:2025cns ; Caldwell:2025inn . In addition, NEC violation also provides a compelling mechanism for the formation of PBHs as a dark matter candidate Cai:2023uhc , as well as for generating stochastic gravitational-wave background signals detectable by pulsar timing array observations Zhu:2023lbf ; Ye:2023tpz , see also Cai:2020qpu ; Cai:2022nqv ; Cai:2022lec ; Pan:2024ydt ; Li:2024oru ; Li:2025ilc ; Lai:2025xov ; Lai:2025efh .
Although widely employed in theoretical models, the viability of NEC violation in the real world remains in question. Under the perfect-fluid assumption, NEC violation means the total energy density plus pressure becomes negative along a null direction, which allows exotic behaviors like an increase of energy density even in an adiabatically expanding universe. For theorists, NEC violation is often viewed unfavorably because it tends to trigger pathological instabilities, including ghost and gradient instabilities Libanov:2016kfc ; Kobayashi:2016xpl , see also Battarra:2014tga ; Koehn:2015vvy ; Ijjas:2016tpn . It is demonstrated in the context of nonsingular cosmology that fully stable NEC violation can be constructed in “beyond Horndeski” theories Cai:2016thi ; Creminelli:2016zwa ; Cai:2017tku ; Cai:2017dyi ; Kolevatov:2017voe . Furthermore, the NEC can notoriously be violated at the quantum level, for instance in the quantum Casimir effect Brown:1969na and in the Hawking evaporation process (see, e.g. Kontou:2023ntd for a recent update on this topic). Notably, in the cosmological context, NEC violation can also emerge from the secular growth of vacuum expectation values for self-interacting scalar fields on a de Sitter background, as derived within the framework of perturbative quantum field theory Kahya:2009sz .
@@ II.2 The SNEC conjecture
where $t_{\mathrm{s}}$ and $t_{\mathrm{e}}$ denote the beginning and the end of the cosmic interval we are interested in, respectively. Here, we have already used $\mathrm{d}\lambda=a\mathrm{d}t$. It is interesting to note that the secular growth of $\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle\approx-\frac{4}{3}\frac{\lambda H^{4}}{(4\pi)^{4}}\ln a(t)$ for a massless, minimally coupled scalar with quartic self-interaction on a non-dynamical de Sitter background (as computed in Kahya:2009sz with appropriate initial state corrections) respects the SNEC, even though it violates the classical NEC (see the Appendix. A for details). In the following, (10) serves as the master condition for examining the Genesis scenario.
@@ Acknowledgements.
We thank Yun-Song Piao for valuable discussions. Dong-Hui Yu and Yong Cai are supported in part by the National Natural Science Foundation of China (Grant Nos. 12575066, 11905224) and the Natural Science Foundation of Henan Province (Grant No. 242300420231). Mian Zhu is supported by the National Natural Science Foundation of China (Grant No. 12503005), and the Fundamental Research Funds for the Central Universities Grant No. YJ202551. The data that support the findings of this article are openly available dataYu260228 .
Appendix A Consistency of SNEC with perturbative secular growth in de Sitter space
The secular growth of $\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle\propto-\frac{4}{3}\frac{\lambda H^{4}}{(4\pi)^{4}}\ln[a(t)]$ for a massless, minimally coupled scalar with quartic self-interaction on non-dynamical de Sitter background (as computed in Kahya:2009sz with appropriate initial state corrections) is well-founded and deserves careful examination. In the following, we check whether this example satisfies the SNEC.
A.1 A geometric perspective on the SNEC
The energy condition is of vital importance in cosmology as it connects the physics of spacetime and the matter content. In the context of primordial universe, we are mainly interested in the property of geometry instead of matter content. For instance, the famous “no-go” theorem in Ref. Borde:1993xh ; Borde:2001nh concerns on the geodesic completeness of the spacetime, i.e., whether the geodesics of a particle terminates at a finite past. The proof is purely geometric, as only the FLRW metric is involved. Thus, their result actually states that an FLRW universe satisfying Null Convergence Condition (NCC) must encounter the initial singularity. Till this step, there is nothing to do with the energy condition of matter content. Now working in general relativity, the Einstein equation relates $G_{\mu\nu}$ and $T_{\mu\nu}$, and the violation of NCC is equivalent to NEC violation, and the statement that “the free of initial singularity requires NEC violation” is made in this sense.
The similar logic applies to the study of SNEC. In Ref. Freivogel:2018gxj , the foundation paper of SNEC, the authors state that “The quantity that is bounded by this conjecture is related by the Raychadauri equation to the focusing of null geodesics. As we motivate in more detail below, this conjecture claims that whenever geodesics ‘de-focus’ due to negative null energy, they do so by a small amount that remains in the linear gravity regime.” (which can be found in the paragraphs below Eq. (4) of that paper). Namely, the SNEC bound is motivated by the property of null geodesics, firstly written as $R_{kk}^{s}\geq-\#/\tau^{2}$ indicated in Eq. (6) in that paper, and then translated into $T_{kk}^{s}\geq-\#/\tau^{2}$ using Einstein equation.
In the setup of Ref. Kahya:2009sz , the geometry is assumed to be a non-dynamical de Sitter background, which implies that $H$ is a constant. Consequently, it seems that both the NCC (i.e., $R_{\mu\nu}k^{\mu}k^{\nu}\geq 0$) and the SNEC are satisfied by the scenario studied in Ref. Kahya:2009sz , since $R_{\mu\nu}k^{\mu}k^{\nu}=-2\dot{H}(k^{0})^{2}=0$ for the spatially flat Friedmann universe.
A.2 A matter perspective and the backreaction
For a massless, minimally coupled scalar with quartic self-interaction on a fixed non-dynamical de Sitter background, after applying the initial state corrections proposed in Kahya:2009sz , the expectation value is
$\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle\propto\frac{\lambda H^{4}}{(4\pi)^{4}}\times\left(-\frac{4}{3}\right)\ln a+O(\lambda^{2})\,$, (47)
which can be derived from Eqs. (3) and (4) of Ref. Kahya:2009sz . This expression indeed violates the classical NEC since $\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle<0$ for $a(t)>1$. As the universe expands, the accumulated negative energy grows logarithmically with the scale factor ($\propto\ln a$) and it appears sufficient to violate the SNEC eventually. However, in a self-consistent theory of semiclassical gravity, one must consider the backreaction of these quantum fluctuations on the spacetime geometry. We demonstrate below that the backreaction becomes non-negligible well before the accumulated negative energy is sufficient to violate the SNEC.
To make this argument quantitative, we compare the timescale (or $e$-folding number) required to violate SNEC ($N_{\text{SNEC}}$) against the timescale at which the background geometry significantly deviates from de Sitter space ($N_{\text{BR}}$).
A.2.1 The Timescale for SNEC Violation
First, we derive the condition for SNEC violation. Consider a flat FLRW metric $\mathrm{d}s^{2}=-\mathrm{d}t^{2}+a^{2}(t)\mathrm{d}\vec{x}^{2}$. Let $k^{\mu}=(k^{0},k^{i})$ be a null vector tangent to the geodesic $\gamma(\lambda)$ with affine parameter $\lambda$, where $k^{0}=a|\vec{k}|$ and $\mathrm{d}t/\mathrm{d}\lambda=k^{0}$. We have $\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle=(\rho+p)(k^{0})^{2}$. According to Eqs. (3) and (4) of Kahya:2009sz , for a $\lambda\phi^{4}$ theory,
$|(\rho+p)_{Q}|\approx\frac{\lambda H^{4}}{(4\pi)^{4}}\left(\frac{4}{3}N\right)\,$, (48)
where $N=\ln a$ is the $e$-folding number, the subscript “Q” denotes contributions from quantum corrections.
The SNEC requires $\int\mathrm{d}\lambda f^{2}(\lambda)\langle T_{\mu\nu}k^{\mu}k^{\nu}\rangle\geq-B/(G\sigma^{2})$, where $\sigma$ is the smearing scale in affine units. Since $\ln a$ varies slowly over a Hubble time ($\Delta N\sim 1\ll N$ at late times), we can approximate the integral by the local value. To obtain the most physically relevant constraint, the physical smearing time $\Delta t$ is chosen to be of the order of the curvature scale, $\Delta t\sim H^{-1}$ (see e.g. Ref. Moghtaderi:2025cns ). The corresponding affine smearing scale is $\sigma\approx\Delta\lambda=\Delta t/k^{0}\approx(k^{0}H)^{-1}$. Substituting these into the SNEC inequality, we find
$(\rho+p)(k^{0})^{2}\gtrsim-\frac{B}{G[(k^{0}H)^{-1}]^{2}}=-\frac{B(k^{0})^{2}H^{2}}{G}\,.$ (49)
We observe that the factor $(k^{0})^{2}$ cancels out on both sides, yielding a bound:
$|(\rho+p)|\lesssim 8\pi BM_{\mathrm{P}}^{2}H^{2}\,.$ (50)
Note that we have used $8\pi G=M_{\mathrm{P}}^{-2}$. Equating the quantum source term to this bound, we find the timescale $N_{\text{SNEC}}$:
$\frac{\lambda H^{4}}{(4\pi)^{4}}\frac{4}{3}N_{\text{SNEC}}\approx 8\pi BM_{\mathrm{P}}^{2}H^{2}\quad\Longrightarrow\quad N_{\text{SNEC}}\approx\frac{6\pi B(4\pi)^{4}}{\lambda}\left(\frac{M_{\mathrm{P}}}{H}\right)^{2}\,.$ (51)
Therefore, the SNEC will be violated when the universe expands to the scale where $\ln a=N_{\text{SNEC}}$.
A.2.2 Comparison with Backreaction Timescales
We now evaluate when the fixed background approximation breaks down. We consider two distinct criteria for the onset of significant backreaction.
Criterion I: Geometric Deviation (Slow-roll). We can define backreaction based on the induced time-evolution of the Hubble parameter $H$. Using $2M_{\mathrm{P}}^{2}|\dot{H}|=|(\rho+p)|$, we define $N_{\text{BR1}}$ at the time when the parameter $\mathcal{E}\equiv|\dot{H}/H^{2}|$ reaches a threshold $\mathcal{E}_{0}$, i.e.,
$\frac{\lambda H^{4}}{(4\pi)^{4}}\frac{4}{3}N_{\text{BR1}}=2\mathcal{E}_{0}M_{\mathrm{P}}^{2}H^{2}\quad\Longrightarrow\quad N_{\text{BR1}}\approx\frac{3\mathcal{E}_{0}(4\pi)^{4}}{2\lambda}\left(\frac{M_{\mathrm{P}}}{H}\right)^{2}\,.$ (52)
The ratio of timescales is
$\frac{N_{\text{SNEC}}}{N_{\text{BR1}}}=\frac{4\pi B}{\mathcal{E}_{0}}\,.$ (53)
For typical values of those parameters (e.g., $B\approx 1/32\pi$ and a conservative $\mathcal{E}_{0}\sim 10^{-1}$), this ratio is greater than unity ($N_{\text{SNEC}}\gtrsim N_{\text{BR1}}$). This indicates that when the SNEC is violated, the evolution of $H$ over time has become sufficiently significant such that the fixed background approximation is nearly broken down.
Criterion II: Energy Density Dominance. The leading order contribution to the energy density grows quadratically with $N$. From Eq. (3) of Kahya:2009sz , $\rho_{Q}\approx\frac{\lambda H^{4}}{(4\pi)^{4}}2N^{2}$. We define the backreaction time $N_{\text{BR2}}$ as the moment when the quantum energy density reaches a fraction $\varepsilon$ of the background energy density $3M_{\mathrm{P}}^{2}H^{2}$, i.e.,
$\frac{\lambda H^{4}}{(4\pi)^{4}}2N_{\text{BR2}}^{2}=\varepsilon\cdot 3M_{\mathrm{P}}^{2}H^{2}\quad\Longrightarrow\quad N_{\text{BR2}}\approx\sqrt{\frac{3\varepsilon(4\pi)^{4}}{2\lambda}}\left(\frac{M_{\mathrm{P}}}{H}\right)\,.$ (54)
Comparing the two timescales, we find
$\frac{N_{\text{SNEC}}}{N_{\text{BR2}}}\approx\pi B\sqrt{\frac{72}{\varepsilon\lambda}}\left(\frac{M_{\mathrm{P}}}{H}\right)\gg 1\,.$ (55)
Since $M_{\mathrm{P}}/H$ is typically very large (e.g., $10^{5}$), $B\approx 1/32\pi$, and $\rho_{Q}$ grows as $N^{2}$ while $|(\rho+p)_{Q}|$ grows only as $N$, the backreaction from the energy density will destroy the de Sitter background long before the SNEC bound is challenged.
Therefore, the SNEC is satisfied by the example given by Ref. Kahya:2009sz from both a geometric perspective and a matter perspective. Intriguingly, the backreaction becomes significant before the SNEC is violated. As has been pointed out in the footnote on page 3 of Kahya:2009sz , “these secular corrections eventually become nonperturbatively strong”. The result given by Eq. (47) relies on the assumption of a background that remains fixed despite the growth in the stress-energy tensor. In a physically realistic universe, the geometry would evolve (breaking the de Sitter symmetry) well before the secular growth of the NEC-violating term could break the SNEC bound. Therefore, the SNEC is a consistent constraint for dynamical cosmological models.
Additionally, it is also interesting to consider the phenomenon of “self-reproduction” during inflation Linde:1986fd . While quantum fluctuations can indeed drive the scalar field up its potential in certain regions (increasing local $H$), they simultaneously drive the field down in other regions (decreasing $H$). The SNEC is a constraint on the semi-classical expectation value $\langle T_{\mu\nu}\rangle$ along a geodesic. Those localized fluctuations do not imply that the integrated expectation value $\int\langle T_{\mu\nu}\rangle k^{\mu}k^{\nu}f(\lambda)d\lambda$ along a geodesic becomes negatively divergent. The ensemble average of the energy-momentum tensor remains controlled by the competition between classical drift and quantum diffusion. Thus, a local, stochastic increase in $H$ in a “self-reproducing” patch does not automatically imply that the semi-classical integral condition is violated for the entire spacetime or for the Genesis scenario we are investigating.